OpenStudy (babyslapmafro):

Please help: Convert the following cartesian coordinate into polar coordinate. (-sqrt2,sqrt2)

OpenStudy (babyslapmafro):

these are the possible answers: 2, pi/4 2, 3pi/4 4, pi/4 4, 3pi/4

OpenStudy (babyslapmafro):

i thought it was 3pi/4,7pi/4

OpenStudy (freckles):

so you have (x,y) you want to find (r,theta) \[r=\sqrt{x^2+y^2}\] and since our point is in the second quadrant to find theta you do \[\theta=\arctan(\frac{y}{x})+\pi\]

OpenStudy (freckles):

 you can also put +2pi to follow any of those for the theta part since going a full rotation after that point takes you back to that point